In a titration experiment, 29.9 mL of 0.797 M HCOOH neutralizes 20.6 mL of Ba(OH)2. What is the concentration of the Ba(OH)2 solution?
Answer #1

The balanced equation for the reaction between HCOOH and Ba(OH)2 is

2 HCOOH + Ba(OH)2 ----------> Ba(HCOO)2 + 2 H2O

Number of moles of HCOOH = M*V = 0.797 M * 29.9 mL = 23.8303 mmol

According to balanced equation,

2 moles of HCOOH reacts with 1 mole of Ba(OH)2

So, 23.8303 mmol of HCOOH would react with 23.8303 mmol * ½ = 11.91515 mmol of Ba(OH)2

Number of moles ofBa(OH)2 = 11.91515 mmol

Volume of Ba(OH)2 solution = moles/molarity = 11.91515 mmol/20.6 mL = 0.578405 M

Volume of Ba(OH)2 = 0.578 M